# Designing Concise ConvNets with Columnar Stages

### v1

@@ 4.2 Comparison with Standard ConvNets
CoSNet vs recent VanillaNet Chen et al. (2024). CoSNet rivals recent ConvNet design, VanillaNet. VanillaNet is shallow and mainly focuses on latency. Our CoSNet-A00 shows similar latency at fewer parameters, fewer FLOPs, and high accuracy compared with VanillaNet-$6$ (Table 1).
CoSNet vs recent ParNet Goyal et al. (2021). CoSNet outperforms recent non-deep ParNet that focuses on lower latency (Table 1 R4). CoSNet is uni-branched, while ParNet has multiple shallow branches which serialize the computations, thus making them deeper virtually.
CoSNet vs RepVGG Ding et al. (2021). RepVGG offers a plain VGG-like Simonyan & Zisserman (2014) structure via Structural Reparameterization (SR). However, its training complexity is high due to a large number of parameters and three branches at each layer (Figure 1f). Hence, we show its performance with and without SR.
@@
CoSNet vs EfficientNet Tan & Le (2019). Although we do not aim for a mobile regime in this paper, we show that having fewer parameters and FLOPs does not guarantee faster speeds. As shown in Table 1, EfficientNet-B00 has $50\%$ fewer parameters and $77\%$ fewer FLOPs, but is $50\%$ deeper, and runs $37\%$ slower. By exploring the design space, CoSNet can be extended to the mobile regime.
CoSNet vs ResNet He et al. (2016) family. As shown in Table 1, CoSNet-A00 is $6\%$ more accurate, has $25\%$ fewer parameters, shows similar runtime, and shows $31\%$ fewer FLOPs than ResNet-$18$ although CoSNet has $6$ more layers. Similarly, in contrast to ResNet-$34$, it is more accurate by $3\%$ with $59\%$ fewer parameters, $66\%$ fewer FLOPs, and $23\%$ less layers, while it is fast by $37\%$. ResNet-$50$ is the widely employed backbone in downstream tasks Carion et al. (2020); Goyal et al. (2017); He et al. (2017); Ren et al. (2015) due to its affordability regarding representation power, FLOPs, depth, and accuracy. Table 1 shows that CoSNet-B00 surpasses ResNet-$50$ while being $50\%$ shallower, $22\%$ fewer parameters, $25\%$ fewer FLOPs, and $40\%$ faster.
CoSNet vs bigger ResNet He et al. (2016) and ResNeXt Xie et al. (2017) models. As shown in Table 1, CoSNet-C1111 is better than bigger variants of ResNet, which serves as backbones for cutting-edge works Carion et al. (2020); Li et al. (2022). Our CoSNet outperforms them in various aspects while being $72\%$ and $82\%$ less deep relative to ResNet-$101$ and ResNet-$152$, respectively. CoSNet also runs faster by $50\%$ in $50\%$ fewer parameters and FLOPs. In addition, despite being smaller than ResNeXt Xie et al. (2017), CoSNet-C1111 outperforms it in various aspects. Overall CoSNet-C1111 is $50\%$ less deeper than ResNeXt-$50$ while running $50\%$ faster at $6\%$ fewer FLOPs, $2\%$ fewer parameters while being more accurate. In contrast to ResNeXt-$101$, CoSNet-C2222 is $75\%$ less deeper, $11\%$ fewer parameters, $12\%$ fewer FLOPs, and $35\%$ faster at a higher accuracy.
@@ 4.3 Additional Experiments
CoSNet has small training walltime. We provide an additional comparison with the recent ConvNet design, VanillaNet Chen et al. (2024), under training settings. Table 2 shows that despite VanillaNet being a shallow network, it has a high training time. We speculate that the large number of channels in the deeper layers of VanillaNet slows down batch processing at large batch sizes. In CoSNet, parallel columnar convolutions and controlled parameter growth in the deeper layers counter this issue, leading to lower training time.
CoSNet is seamlessly compatible with SE-like Hu et al. (2018) modules. Table 3 shows the results when CoSNet is used in conjunction with Squeeze and Excitation (SE) like modules Hu et al. (2018). It outperforms recent attention mechanism (AFF Dai et al. (2021a), SKNet Li et al. (2019), and CBAM Woo et al. (2018)) applied to ResNet-50.
@@ 4.4 CoSNet in State-of-the-art Detection Transformer
We apply CoSNet to state-of-the-art object Detection Transformer, DN-DETR Li et al. (2022) to demonstrate the effectiveness of CoSNet in the downstream task. We experiment on MS-COCO Lin et al. (2014) benchmark and utilized DN-DETR’s default training settings.
@@ 4.5 Visualization of Attention
To comprehend CoSNet’s better performance, we investigate its class activation maps (CAM) on ImageNet Deng et al. (2009) validation set. We use CAM output from popular Full-Grad-CAM Srinivas & Fleuret (2019) for a given class. CAM visualizations of ResNet-$50$ and CoSNet-B1111 are shown in Figure 6. It can be seen that CoSNet, despite being $50\%$ shallower than ResNet, is better at learning to attend regions of the target class relative to the baseline.
@@ 5 Conclusion
Limitations. CoSNet is open for improvement. In this paper, we have built a simple template architecture that can further evolve like ConvNext Liu et al. (2022). For instance, a comprehensive design space of CoSNet including mobile regime can be explored, similar to RegNet Radosavovic et al. (2020). Besides, layer merging post-training, shown in VanillarNet Chen et al. (2024), can be utilized to develop more shallower variants of CoSNet. In addition to that, CoSNet can also be married with a Transformer attention mechanism like Dai et al. (2021b) or Liu et al. (2023).
@@ Appendix B Ablation Study
Appendix C Training setting
We train models in PyTorch pytorch using eight NVIDIA A40 GPUs.
Appendix D PyTorch Code
All codes shall be open-sourced in PyTorch pytorch post the review process. Here, we provide a code snippet of a CoSNet-Unit. Please see until the end of this document.
Appendix E Complete Network Visualization

### v2

@@ 4.2 Comparison with Standard ConvNets
CoSNet vs recent VanillaNet (Chen et al., 2023). CoSNet rivals recent ConvNet design, VanillaNet. VanillaNet is shallow and mainly focuses on latency. Our CoSNet-A00 shows similar latency at fewer parameters, fewer FLOPs, and high accuracy compared with VanillaNet-$6$ (Table 1).
CoSNet vs recent ParNet (Goyal et al., 2021). CoSNet outperforms recent non-deep ParNet that focuses on lower latency (Table 1 R4). CoSNet is uni-branched, while ParNet has multiple shallow branches which serialize the computations, thus making them deeper virtually.
CoSNet vs RepVGG (Ding et al., 2021). RepVGG offers a plain VGG-like (Simonyan & Zisserman, 2014) structure via Structural Reparameterization (SR). However, its training complexity is high due to a large number of parameters and three branches at each layer (Figure 1f). Hence, we show its performance with and without SR.
@@
CoSNet vs EfficientNet (Tan & Le, 2019). Although we do not aim for a mobile regime in this paper, we show that having fewer parameters and FLOPs does not guarantee faster speeds. As shown in Table 1, EfficientNet-B00 has $50\%$ fewer parameters and $77\%$ fewer FLOPs, but is $50\%$ deeper, and runs $37\%$ slower. By exploring the design space, CoSNet can be extended to the mobile regime.
CoSNet vs ResNet (He et al., 2016) family. As shown in Table 1, CoSNet-A00 is $6\%$ more accurate, has $25\%$ fewer parameters, shows similar runtime, and shows $31\%$ fewer FLOPs than ResNet-$18$ although CoSNet has $6$ more layers. Similarly, in contrast to ResNet-$34$, it is more accurate by $3\%$ with $59\%$ fewer parameters, $66\%$ fewer FLOPs, and $23\%$ less layers, while it is fast by $37\%$. ResNet-$50$ is the widely employed backbone in downstream tasks (Carion et al., 2020; Goyal et al., 2017; He et al., 2017; Ren et al., 2015) due to its affordability regarding representation power, FLOPs, depth, and accuracy. Table 1 shows that CoSNet-B00 surpasses ResNet-$50$ while being $50\%$ shallower, $22\%$ fewer parameters, $25\%$ fewer FLOPs, and $40\%$ faster.
CoSNet vs bigger ResNet (He et al., 2016) and ResNeXt (Xie et al., 2017) models. As shown in Table 1, CoSNet-C1111 is better than bigger variants of ResNet, which serves as backbones for cutting-edge works (Carion et al., 2020; Li et al., 2022). Our CoSNet outperforms them in various aspects while being $72\%$ and $82\%$ less deep relative to ResNet-$101$ and ResNet-$152$, respectively. CoSNet also runs faster by $50\%$ in $50\%$ fewer parameters and FLOPs. In addition, despite being smaller than ResNeXt (Xie et al., 2017), CoSNet-C1111 outperforms it in various aspects. Overall CoSNet-C1111 is $50\%$ less deeper than ResNeXt-$50$ while running $50\%$ faster at $6\%$ fewer FLOPs, $2\%$ fewer parameters while being more accurate. In contrast to ResNeXt-$101$, CoSNet-C2222 is $75\%$ less deeper, $11\%$ fewer parameters, $12\%$ fewer FLOPs, and $35\%$ faster at a higher accuracy.
@@ 4.3 Additional Experiments
CoSNet has small training walltime. We provide an additional comparison with the recent ConvNet design, VanillaNet (Chen et al., 2023), under training settings. Table 2 shows that despite VanillaNet being a shallow network, it has a high training time. We speculate that the large number of channels in the deeper layers of VanillaNet slows down batch processing at large batch sizes. In CoSNet, parallel columnar convolutions and controlled parameter growth in the deeper layers counter this issue, leading to lower training time.
CoSNet is seamlessly compatible with SE-like (Hu et al., 2018) modules. Table 3 shows the results when CoSNet is used in conjunction with Squeeze and Excitation (SE) like modules (Hu et al., 2018). It outperforms recent attention mechanism (AFF (Dai et al., 2021a), SKNet (Li et al., 2019), and CBAM (Woo et al., 2018)) applied to ResNet-50.
@@ 4.4 CoSNet in State-of-the-art Detection Transformer
We apply CoSNet to state-of-the-art object Detection Transformer, DN-DETR (Li et al., 2022) to demonstrate the effectiveness of CoSNet in the downstream task. We experiment on MS-COCO (Lin et al., 2014) benchmark and utilized DN-DETR’s default training settings.
@@ 4.5 Visualization of Attention
To comprehend CoSNet’s better performance, we investigate its class activation maps (CAM) on ImageNet (Deng et al., 2009) validation set. We use CAM output from popular Full-Grad-CAM (Srinivas & Fleuret, 2019) for a given class. CAM visualizations of ResNet-$50$ and CoSNet-B1111 are shown in Figure 6. It can be seen that CoSNet, despite being $50\%$ shallower than ResNet, is better at learning to attend regions of the target class relative to the baseline.
@@ 5 Conclusion
Future work. CoSNet is open for improvement. In this paper, we have built a simple template architecture that can further evolve like ConvNext (Liu et al., 2022). For instance, a comprehensive design space of CoSNet including mobile regime can be explored, similar to RegNet (Radosavovic et al., 2020). Besides, layer merging post-training, shown in VanillarNet (Chen et al., 2023), can be utilized to develop shallower variants of CoSNet. In addition to that, CoSNet can also be married with a Transformer attention mechanism like (Dai et al., 2021b) or (Liu et al., 2023).
Acknowledgements. Jaesik Park was supported by MSIT grant (RS-2021-II211343: AI Graduate School Program at Seoul National University (5%) and 2023R1A1C200781211 (95%))
@@ Appendix B Ablation Study
Appendix C The effect of batch sizes of the baseline approaches.
In the literature, some baselines are trained with larger batch sizes (above 1024), but others have been trained at a much smaller batch size ($256$). Therefore, we retrained high batch size baselines with $256$ batch sizes to avoid getting biased conclusions about the effects of large batch sizes. Such results with 256 batch size are carefully reported in Table 1.
In this section, we present the results of the baselines with larger batch sizes in Table A3. As widely studied, the baseline approaches Tan & Le (2019); Liu et al. (2023; 2022) show improved accuracy. Interestingly, it can be noticed that CoSNet trained with a 256 batch size can compete with state-of-the-art approaches trained with a larger batch size. This shows the utility of obtaining higher accuracies in resource-constrained training scenarios (i.e., limited memory to fit 4096 batch, etc.).
Table A3: Effect of batch size on the baselines and CoSNet in the context.
[…]
$\bullet$ EfficientNet-B00 Tan & Le (2019) ConvNet $256$ $49$ $5.3$M $0.40$B $8$ms $125$ $75.1$
$\bullet$ EfficientNet-B00 Tan & Le (2019) ConvNet $2048$ $49$ $5.3$M $0.40$B $8$ms $125$ $77.1$
$\bullet$ EfficientViT-M5555 Liu et al. (2023) Transformer $256$ $70$ $12.4$M $0.60$B $7$ms $142$ $76.8$
$\bullet$ EfficientViT-M5555 Liu et al. (2023) Transformer $2048$ $70$ $12.4$M $0.60$B $7$ms $142$ $77.1$
$\bullet$ CoSNet-A00 ConvNet $256$ $26$ $8.8$M $1.25$B $6$ms $167$ $77.1$
$\bullet$ CoSNet-A1111-PFF ConvNet $256$ $38$ $12.7$M $1.93$B $7$ms $143$ $79.7$
$\bullet$ ConvNeXt-T Liu et al. (2022) ConvNet $256$ $59$ $29.0$M $4.50$B $13$ms $77$ $81.8$
$\bullet$ ConvNeXt-T Liu et al. (2022) ConvNet $4096$ $59$ $29.0$M $4.50$B $13$ms $77$ $82.1$
$\bullet$ CoSNet-C2222 ConvNet $256$ $26$ $38.9$M $7.09$B $11$ms $90$ $82.1$
$\bullet$ CoSNet-B2222-PFF ConvNet $256$ $38$ $34.3$M $5.91$B $10$ms $100$ $82.7$
Table A4: CoSNet in RetinaNet $x1$ Lin et al. (2017) object detection framework.
Method #Depth #Params AP APS APM APL
$\bullet$ EfficientViT-M4444 Liu et al. (2023) $42$ $8.8$M $32.7$ $17.6$ $35.3$ $46.0$
[…]
Table A5: CoSNet in PSPNet Zhao et al. (2017) semantic segmentation framework.
Method #Params mIoU FPS
$\bullet$ RepVGG-B1111g2222 Ding et al. (2021) $41.36$M $78.88$ $13$
$\bullet$ ResNet-$50$ $25.5$M $77.17$ $13$
$\bullet$ CosNet-B1 $22.0$M $79.05$ $17$
Appendix D Additional Results
Table A4 shows results on RetinaNet $x1$ Lin et al. (2017) detection pipeline. It can be seen that, for a comparable vision transformer backbone, CoSNet performs better. We also provide semantic segmentation results for the popular PSPNet pspnet semantic segmentation framework. It can be seen that CoSNet performs better than the baselines.
Appendix E Training setting
We train models in PyTorch Paszke et al. (2019) using eight NVIDIA A40 GPUs.
Appendix F Complete Network Visualization